# The unreasonable likelihood of being Origin of life, terraforming, and AI

### v1

@@ Abstract
“All cells come from cells” [1] leads us into a classic chicken-and-egg dilemma: where did the first cell come from? Either it came from somewhere else—conveniently outsourcing the mystery—or it emerged from the laws of physics and chemistry on a young, chaotic, and geologically active Earth. Once a minimal protocell or replicator emerged, Darwinian evolution could take over, merely selecting and diversifying its way toward greater complexity, biodiversity, and eventually, the modern biosphere, from gold fish to particle accelerators and social media. One thing this onset of a discussion illustrates - how little we know about our origin, and the cosmos in general.
@@
As for the “somewhere else” hypothesis, science fiction has long entertained terraforming scenarios. In 2001: A Space Odyssey, an inscrutable black monolith gives early hominins a mysterious push toward cognition—and tool use [2]. In Star Trek, entire planetary ecosystems are engineered by advanced civilizations or errant AI, often with remarkable speed [3]. The X-Files proposed that alien viruses might have seeded Earth, albeit oddly fixated on rural America [4]. In Prometheus, a lone extraterrestrial—credited only as “the Engineer”—dissolves himself in a waterfall, presumably catalyzing Earth’s biochemistry, though the narrative coherence dissolves nearly as fast [5]. The mother of all terraforming stories is, however, the Book of Genesis, written more than 2,500 years ago. It declares that life was created in six days [6] — a claim that has profoundly shaped human thought, even if not amenable to scientific testing.
@@
One step closer to science (though not necessarily to certainty), Francis Crick—the co-discoverer of DNA’s structure—and Leslie Orgel proposed in 1973 the idea of directed panspermia [7]. In their scenario, an advanced extraterrestrial civilization, facing extinction or perhaps scientific curiosity, dispatches microbial “starter kits” to habitable planets like ours (see Fig. 1 for an illustration). While Crick and Orgel attempted to formulate this idea more like a testable hypothesis, it deftly relocates the explanatory burden to someone else’s biochemistry.
@@
Meanwhile, modern Bayesian analyses of life’s emergence struggle with deeply unconstrained data (or lack thereof): a sample size of one, and priors shaped not only by empirical observation but also by immense subjectivity [8, 9]. There’s also the problem of timing—life appeared early on Earth, yet the intelligent observer required to marvel this fact arrived very much later, introducing a strong observational bias. In this work, we revisit the origin of life through the lens of modern cell models and AI-based tools, to better appreciate the informational and physical complexity required for protocell formation. We begin with general considerations and propose a qualitative theoretical framework. Finally, we follow Feynman’s mantra that ”doubt and uncertainty are not to be feared but welcomed” [10], and attempt to cautiously conclude something.
@@
When did life emerge, and what are the soft and hard constraints? There were two massive impacts with global sterilizing effects: Earth was first struck by a Mars-sized body, named “Theia,” about 4.51 Gy ago, leading to the formation of the Moon [11]. This was followed by another sterilizing impact, dubbed “Moneta,” approximately 40 My later [12]. Mineralogical evidence from zircons indicates that both an atmosphere and liquid water were already present on Earth’s surface around 4.404 Gy ago [13]. The oldest putative evidence for life comes from 13C-depleted carbon inclusions in 4.1 Gy-old zircon deposits [14]. The oldest direct and undisputed evidence is found in microfossils preserved in 3.465 Gy-old rocks in Western Australia [15, 16, 17].
@@
Recently, a study combining molecular clock methodology, horizontal gene-transfer-aware phylogenetic reconciliation, and existing biogeochemical models addressed the nature of the last universal common ancestor (LUCA), estimating its age, gene content, metabolism, and ecological context [18]. Surprisingly, LUCA is estimated to have lived around $\sim 4.2$ Gy ago, based on calibration with fossil and isotope records. More perplexing still, LUCA appears to have been an anaerobic acetogen, metabolically similar to modern prokaryotes, and already equipped with ATP synthesis, a TCA cycle, early immune capabilities including CRISPR-Cas effector proteins, and embedded within a microbial ecosystem. A recent consensus across eight genomic and proteomic studies broadly supports this view [19]. This suggests that a protocell—presumably far simpler—must have emerged even earlier, perhaps relying on RNA or other primitive replicator systems.
@@
Only in the last blink of planetary time — some 100,000 years ago — did anatomically modern humans appear, wielding curiosity and stunning technological inventions. Although the controlled use of fire by earlier hominins dates back about 1 Myr [20], in geological terms intelligence has arrived extraordinarily late. Yet anthropic considerations must be taken into account: without human observers, it would not have been possible to ponder this spectacle. Intriguingly, this late arrival may be further constrained from the other end of the timeline, as Earth’s biosphere is projected to collapse in roughly 100 Myr due to impending “CO2 starvation” (even way before the sun becomes a Red Giant and engulfs the Earth) [21]. As solar luminosity increases, silicate weathering accelerates, drawing down atmospheric CO2. Below $\sim 150$ ppm, photosynthesis ceases, triggering ecological collapse. Hence, if intelligent life requires too much time to emerge, many planets—especially around hotter, short-lived stars—may simply run out of time, biasing observations toward an unusually early start.
@@
On a more optimistic note, a modified Drake equation—updated with exoplanet data—places a firm lower bound on the likelihood that at least one other technological species has arisen at some point in the observable universe. If the probability that a habitable-zone planet spawns a technological species is $\geq 10^{-24}$ — about the odds of winning the lottery a modest $8$ times in a row — then humanity is almost certainly not alone [22]. So statistically speaking, we’re probably not the only ones wondering whether we’re alone.
@@
Since Erwin Schrödinger’s landmark 1944 essay What is Life?, physicists have speculated whether the emergence of life might require physical principles beyond those already known [23]. Schrödinger speculated—presciently, though vaguely—about undiscovered laws governing order and organization in living systems, prompting generations of researchers to explore the physics of self-organization. Since then, the field has progressed dramatically. Here is a short list without any claim of completeness:
@@
Alan Turing’s work on reaction–diffusion systems introduced a mathematical framework for how spatial patterns, such as those observed in morphogenesis, can spontaneously arise from simple rules [24]. The Landauer–Bennett principle revealed a deep connection between information, entropy, and the thermodynamic cost of computation—key concepts for understanding living systems as information-processing entities [25, 26]. Ilya Prigogine’s theory of dissipative structures demonstrated how local order can emerge in open systems that export entropy [27]. Stuart Kauffman’s concept of autocatalytic sets offered a compelling model for how chemical networks might self-amplify and sustain themselves [28]. The latter two demonstrate nicely that complex systems have new emergent collective features beyond the sum of their parts, not predictable from first principles [29].
More recently, in work bordering the popular science literature, Jeremy England has proposed that life-like order could emerge generically through driven dissipation under physical selection-like conditions [30], while Sara Walker has emphasized the centrality of information flow and causal structure in defining what it means to be alive [31]. Both argue that to understand life’s origin, we must look beyond molecular inventories and instead focus on the dynamics of information, energy, and organization in physical systems. If life emerged through a relay of quasi-stable, far-from-equilibrium states — each more improbable than the last yet somehow energetically favored — then the transition becomes not miraculous, but thermodynamically admissible. The origin of life would simply be extension of the classical rule that ’exergonic reactions occur spontaneously’ [32].
@@
Broadly speaking, origin-of-life relevant research falls within the rich and rapidly evolving domains of active matter, stochastic thermodynamics, and nonequilibrium phase transitions. These fields have significantly expanded our understanding of how systems far from equilibrium can maintain structure and perform work. Yet despite this progress, a fully satisfying mechanism for the spontaneous emergence of a protocell from disordered chemistry remains elusive.
@@
In science, dogmas define domains of accepted knowledge, enabling structured peer review and funding decisions. Yet they can also inhibit new discoveries [33]. For example, the phrase “nothing in biology makes sense except in the light of evolution” [34], though foundational in lab work and bioinformatics, may inadvertently create boundaries that hinder unconventional thinking about life’s origins. Similarly, the dictum “all cells come from cells” [1] tends to obscure the deeper question of how the first cell came to be.
One thing is certain: with the rise of powerful AI tools and mechanistic models, we now have entirely new ways of exploring biological complexity. Tools like AlphaFold for protein folding [35] and comprehensive whole-cell models [36, 37, 38] allow us to estimate the information content of life using Kolmogorov (algorithmic) complexity. This measures the shortest description required to reproduce a system—the more structured or regular a system, the more it can be compressed and the easier it can be made [39]. Here, by combining such complexity estimates with basic rate-distortion theory [40, 41], we aim to shed new light on the plausibility of life’s rapid emergence.
@@
We approach the problem of life’s origin by estimating how plausible it is for chemical processes to spontaneously assemble a sufficiently complex protocell within the available time on early Earth. We assume this window to be about $500$ Myr—interpolated between “optimistic” ($304$ Myr) and “conservative” ($939$ Myr) estimates [9]. We apply a heuristic version of Shannon’s rate-distortion theory [40, 41], which quantifies the minimal information rate $R(D)$ needed to construct a functioning system, given a tolerated distortion level $D$. Here, $D$ characterizes how much loss (error or degradation) can be tolerated in forming a viable protocell. The distortion rate depends on the entropy of the chaotic prebiotic environment, denoted $H_{\text{prebiotic}}$, which represents the information content of the chemical soup in a pond or hydrothermal vent.
@@
In this simplified formulation, we approximate $R(D)$, typically convex and decreasing, by the ratio of prebiotic entropy, $H_{\text{prebiotic}}$, and molecular persistence time, $D$, where $\eta$ is an efficiency prefactor (also called a ”fudge factor”). This term accounts for the fact that not all molecular encounters result in usable contributions. Molecules may be structurally misaligned, chemically unstable, or fail to integrate into a coherent assembly due to the absence of repair mechanisms or spatial constraints in the local environment.
Admittedly, all quantities in Eq. 1—except perhaps $T_{\text{available}}$—are highly uncertain. Nevertheless, following this line of reasoning, the spontaneous origin of life remains plausible if the left-hand side exceeds the right-hand side. If not, we may be missing some crucial ingredient—such as additional evolutionary pathways, unknown physical principles, or external interventions like Crick and Orgel’s directed panspermia (or, more colorfully, aliens terraforming the planet). For the sake of argument, we set aside the ”cheap” statistical fallback that rare events inevitably occur somewhere in a vast universe.
Figure 2: Mathematical framework based on rate-distortion theory. (A) Funnel of collapsing entropy for the emergence of a protocell. (B) Rate-distortion function from information theory, comparing $R(D)\sim H_{\text{prebiotic}}/D$ as a function of $D$ with $R_{\text{min}}$ from observational constraints (see the main text for explanations).
@@
In the absence of a time machine, we estimate conditions on Earth roughly 4 Gy ago. We begin with the left-hand side of Eq. 1, aiming to estimate the entropy of the chemical soup, $H_{\text{prebiotic}}$, and the tolerated error rate $D$—which characterizes how long complex organic molecules persist before degrading. Importantly, the entropy $H_{\text{prebiotic}}$ pertains not to an entire ocean or vent system, but to the volume accessible to a nascent protocell. While the volume of a single protocell may be around 1 femtolitre ($10^{-15}$ l), relevant molecules could diffuse in from a larger surrounding droplet. We seek order-of-magnitude estimates informed by chemical experiments simulating early Earth conditions, meteorite data, and planetary spectroscopy.
@@
In the Miller–Urey experiments, spark discharges in a reducing gas mixture produced a variety of amino acids and other small organic molecules [42]. During the Cassini–Huygens mission, analysis of Titan’s atmosphere revealed that complex photochemistry can generate thousands of distinct organic species, including nitriles, hydrocarbons, and possible prebiotic precursors [43]. Similarly, carbonaceous chondrite meteorites (e.g., Murchison) contain more than 60 amino acids, as well as nucleobases, polyols, carboxylic acids, and related compounds [44]. This is consistent with samples returned from asteroid Bennu by the OSIRIS-REx mission, which revealed a rich inventory of organic compounds—including those preserved in salt-rich brine inclusions in phyllosilicate minerals, dating back to $\sim 4.5$ billion years [45].
@@
Combinatorial chemistry across diverse geochemical settings could plausibly yield thousands of unique organic molecules, suggesting a chemical library on the order of $10^{5}$$10^{6}$ small molecules in the prebiotic environment. Assuming an average information content of $\sim 10$ bits per molecule—representing functional or structural specificity—this allows for roughly $10^{3}$ distinguishable states per molecule (e.g., different side chains, folding propensities, or reactivities). This estimate aligns with protein domain annotations and bits-per-residue values from sequence-function maps [46, 47, 48], as well as compression-based descriptors like SMARTS or InChI [49]. Multiplying, we estimate the prebiotic entropy as $H_{\text{prebiotic}}\sim 10^{6}$$10^{7}$ bits.
@@
Next, we consider the distortion timescale $D$, interpreted as the average lifetime of a complex organic molecule before degradation by hydrolysis, UV damage, oxidation, or further transformation. This defines the “error rate” or persistence window for useful molecular structures. Under plausible early-Earth conditions, complex organics such as ribose and peptides have half-lives ranging from hours to days. For example, ribose degrades with a half-life of $\sim 3$ hours at pH 7 and 37C [50]. Small peptides degrade over hours to days depending on sequence and pH. Nucleobases like adenine are sensitive to solar UV and degrade within days to weeks. Without an ozone layer, degradation could be faster—on the scale of hours to days. Conservatively, we adopt a mean lifetime of $\sim 1$ day, or about $10^{5}$ seconds, for moderately stable molecules in a warm, aqueous, anoxic environment exposed to UV. However, in cryogenic or mineral-protected contexts—such as asteroid interiors—organic molecules may persist for billions of years [45]. The existence of such intergalatic freezers implies that $D$ could vary over many orders of magnitude depending on local geochemical context.
@@
If the molecular lifetime were much shorter—say $D\sim 1$ s, due to rapid diffusion or lack of protective mechanisms—then the required information rate would increase accordingly. The key idea remains: if $R(D)$ is too low to meet $R_{\text{min}}=I_{\text{protocell}}/T_{\text{available}}$, then abiotic emergence becomes implausible without invoking directed evolution, external contributions (e.g., panspermia), or unknown self-organizing principles. The delivery of relatively stable, complex molecules—by meteorites or interstellar dust—could act as a bootstrap for prebiotic evolution [45]. Given the large uncertainties in parameter values, our aim is not to deliver a definitive conclusion but to structure a quantitative discussion.
@@
The right-hand side of Eq. 1 requires an estimate of the complexity of a protocell, $I_{\text{protocell}}$. How can we quantify this complexity, including its structural and dynamic organization? It shouldn’t come as a surprise that cells are not simply bags of well-mixed molecules, but highly structured systems—down to the folding and interaction of their constituent proteins. Moreover, cellular processes are orchestrated in time. If proteins are the musicians, then signaling, metabolic, and genetic pathways are the conductor and the score — guiding timing, harmony, and transitions.
@@
where $I_{g}$ is the genetic information, $I_{s}$ the structural information (e.g. protein folds, cellular organization), and $I_{d}$ the dynamic information encoded in biochemical pathways. Each term is estimated via Kolmogorov (algorithmic) complexity, defined as the length of the shortest program that can realistically generate the relevant component, at least conceptually.

### v2

@@ Abstract
“All cells come from cells” [Virchow1855] leads us into a classic chicken-and-egg dilemma: where did the first cell come from? Either it came from somewhere else—conveniently outsourcing the mystery—or it emerged from the laws of physics and chemistry on a young, chaotic, and geologically active Earth. Once a minimal protocell or replicator emerged, Darwinian evolution could take over, merely selecting and diversifying its way toward greater complexity, biodiversity, and eventually, the modern biosphere, from gold fish to particle accelerators and social media. One thing this onset of a discussion illustrates - how little we know about our origin, and the cosmos in general.
@@
As for the “somewhere else” hypothesis, science fiction has long entertained terraforming scenarios. In 2001: A Space Odyssey, an inscrutable black monolith gives early hominins a mysterious push toward cognition—and tool use [Kubrick1968]. In Star Trek, entire planetary ecosystems are engineered by advanced civilizations or errant AI, often with remarkable speed [Roddenberry1966]. The X-Files proposed that alien viruses might have seeded Earth, albeit oddly fixated on rural America [Carter1993]. In Prometheus, a lone extraterrestrial—credited only as “the Engineer”—dissolves himself in a waterfall, presumably catalyzing Earth’s biochemistry, though the narrative coherence dissolves nearly as fast [Scott2012]. The mother of all terraforming stories is, however, the Book of Genesis, written more than 2,500 years ago. It declares that life was created in six days [GenesisBible] — a claim that has profoundly shaped human thought, even if not amenable to scientific testing.
@@
One step closer to science (though not necessarily to certainty), Francis Crick—the co-discoverer of DNA’s structure—and Leslie Orgel proposed in 1973 the idea of directed panspermia [Crick1973]. In their scenario, an advanced extraterrestrial civilization, facing extinction or perhaps scientific curiosity, dispatches microbial “starter kits” to habitable planets like ours (see Fig. 1 for an illustration). While Crick and Orgel attempted to formulate this idea more like a testable hypothesis, it deftly relocates the explanatory burden to someone else’s biochemistry.
@@
Meanwhile, modern Bayesian analyses of life’s emergence struggle with deeply unconstrained data (or lack thereof): a sample size of one, and priors shaped not only by empirical observation but also by immense subjectivity [Spiegel2012, Kipping2020]. There’s also the problem of timing—life appeared early on Earth, yet the intelligent observer required to marvel this fact arrived very much later, introducing a strong observational bias. In this work, we revisit the origin of life through the lens of modern cell models and AI-based tools to better appreciate the informational and physical complexity required for protocell formation. We begin with general considerations and propose a qualitative theoretical framework. Finally, we follow Feynman’s mantra that ”doubt and uncertainty are not to be feared but welcomed” [Feynman1955], and attempt to cautiously conclude something.
@@
When did life emerge, and what are the soft and hard constraints? There were two massive impacts with global sterilizing effects: Earth was first struck by a Mars-sized body, named “Theia,” about $4.51$ Gy ago, leading to the formation of the Moon [Hartmann1975]. This was followed by another sterilizing impact, dubbed “Moneta,” approximately $40$ My later [Benner2020], and the Late Heavy Bombardment of many smaller impacts between $\sim 4.1$ to $3.8$ Gy, which, in contrast to the earlier giant impacts, are not believed to have been globally sterilizing. Mineralogical evidence from zircons indicates that both an atmosphere and liquid water were already present on Earth’s surface around $4.404$ Gy ago [Wilde2001]. The oldest putative evidence for life comes from 13C-depleted carbon inclusions in $4.1$ Gy-old zircon deposits [Bell2015]. The oldest direct and undisputed evidence is found in microfossils preserved in $3.465$ Gy-old rocks in Western Australia [Schopf2006, Schopf2007, Schopf2018].
Around the same time, the planet’s core was already generating a global magnetic field: palaeomagnetic signatures imply a functioning geodynamo by $\geq 3.5$ Gy as evidenced by permanently magnetized basalt and dacite rocks, and possibly as early as $4.2$ Gy, though such early zircon records remain debated [Tarduno2015HadeanGeodynamo, Olson2013CoreParadox]. This magnetic field likely played an important role in enabling life to establish and persist on the surface of the Earth by shielding it from the damaging solar wind.
@@
Recently, a study combining molecular clock methodology, horizontal gene-transfer-aware phylogenetic reconciliation, and existing biogeochemical models addressed the nature of the last universal common ancestor (LUCA), estimating its age, gene content, metabolism, and ecological context [Moody2024]. Surprisingly, LUCA is estimated to have lived around $\sim 4.2$ Gy ago, based on calibration with fossil and isotope records. More perplexing still, LUCA appears to have been an anaerobic acetogen, metabolically similar to modern prokaryotes, and already equipped with ATP synthesis, a TCA cycle, early immune capabilities including CRISPR-Cas effector proteins, and embedded within a microbial ecosystem. A recent consensus across eight genomic and proteomic studies broadly supports this view [Crapitto2022]. This suggests that a protocell—presumably far simpler—must have emerged even earlier, perhaps relying on RNA or other primitive replicator systems.
@@
Only in the last blink of planetary time — some 100,000 years ago — did anatomically modern humans appear, wielding curiosity and stunning technological inventions. Although the controlled use of fire by earlier hominins dates back about 1 Myr [wrangham2009catching], in geological terms intelligence has arrived extraordinarily late. Yet anthropic considerations must be taken into account: without human observers, it would not have been possible to ponder this spectacle. Intriguingly, this late arrival may be further constrained from the other end of the timeline, as Earth’s biosphere to support human life is projected to collapse in roughly 100 Myr due to impending “CO2 starvation” (even way before the sun becomes a Red Giant and engulfs the Earth) [Lovelock1982Nature, Caldeira1992]. As solar luminosity increases, silicate weathering accelerates, drawing down atmospheric CO2. Below $\sim 150$ ppm, modern plant photosynthesis ceases, triggering ecological collapse  [^1] Photosynthetic pathways differ in their tolerance to low CO2 concentrations. While the biosphere with some cacti and microbes can survive much longer, $\approx 1$ GYrs, this is about $C_{4}$ photosynthesis which can work for even lower CO2 levels (<10<10 ppm CO2). However, most trees, crops, and grass are C3 photosynthesis based, and hence end much earlier (at about 150 ppm CO2). Hence, all animal (and human) life will likely end in about 100 MYrs.. Hence, if intelligent life requires too much time to emerge, many planets—especially around hotter, short-lived stars—may simply run out of time, biasing observations toward an unusually early start.
@@
These considerations highlight that the geological and environmental constraints discussed above are best understood as soft, probabilistic limits rather than hard impossibilities. Such probabilistic limits can produce truly fascinating, even optimistic outlooks: a modified Drake equation—updated with exoplanet data—places a firm lower bound on the likelihood that at least one other technological species has arisen at some point in the observable universe. If the probability that a habitable-zone planet spawns a technological species is $\geq 10^{-24}$ — about the odds of winning the lottery a modest $8$ times in a row — then humanity is almost certainly not alone [Frank2016]. So statistically speaking, we’re probably not the only ones wondering whether we’re alone.
@@
Since Erwin Schrödinger’s landmark 1944 essay What is Life?, physicists have speculated whether the emergence of life might require physical principles beyond those already known [Schrodinger1944]. Schrödinger speculated—presciently, though vaguely—about undiscovered laws governing order and organization in living systems, prompting generations of researchers to explore the physics of self-organization. Since then, the field has progressed dramatically. Here is a short list without any claim of completeness:
@@
Alan Turing’s work on reaction–diffusion systems introduced a mathematical framework for how spatial patterns, such as those observed in morphogenesis, can spontaneously arise from simple rules [Turing1952]. The Landauer–Bennett principle revealed a deep connection between information, entropy, and the thermodynamic cost of computation—key concepts for understanding living systems as information-processing entities [LeffRex2003, Lloyd2006]. Ilya Prigogine’s theory of dissipative structures demonstrated how local order can emerge in open systems that export entropy [Prigogine1984]. Stuart Kauffman’s concept of autocatalytic sets offered a compelling model for how chemical networks might self-amplify and sustain themselves [Kauffman1993]. The latter two demonstrate nicely that complex systems have new emergent collective features beyond the sum of their parts, not predictable from first principles [Anderson1972].
More recently, in work bordering the popular science literature, Jeremy England has proposed that life-like order could emerge generically through driven dissipation under physical selection-like conditions [England2020], while Sara Walker has emphasized the centrality of information flow and causal structure in defining what it means to be alive [Walker2022]. In this spirit, together with Lee Cronin, the emphasis on the history of paths for making complex molecules led to assembly theory (revisited later) [Sharma2023Assembly]. Common to all the above is the argument that to understand life’s origin, we must look beyond molecular inventories and instead focus on the dynamics of information and energy, and thus causation as organizational principles in physical systems. Cast into the language of non-equilibrium physics, the likelihood of reaching a given nonequilibrium state need not be set by the thermodynamic properties of the state itself, but by the entropy produced along the transition path. As shown for multistable non-equilibrium systems, switching with high dissipation can reliably populate even statistically rare states [cook2020thermodynamics]. The origin of life may therefore be reducible to a sequence of exergonic transitions, with path thermodynamics rather than state stability governing the assembly process.
@@
Broadly speaking, origin-of-life relevant research in physics (often theoretical) falls within the rich and rapidly evolving domains of active matter, stochastic thermodynamics, and nonequilibrium phase transitions. These fields have significantly expanded our understanding of how systems far from equilibrium can maintain structure and perform work. Yet despite this progress, a fully satisfying mechanism for the spontaneous emergence of a protocell from disordered chemistry remains elusive. While biological processes obey physical laws, the laws themselves appear agnostic to the emergence of life, given our present understanding.
@@
Beyond formal physical laws, the societal context constrains what is considered worthwhile to work on, as scientific dogmas define domains of accepted knowledge, enabling structured peer review and funding decisions. Yet they can also inhibit new discoveries [Kuhn1962]. For example, the phrase “nothing in biology makes sense except in the light of evolution” [Dobzhansky1973], though foundational in laboratory practice and bioinformatics, may inadvertently create boundaries that hinder unconventional thinking about life’s origins. Similarly, the dictum “all cells come from cells” [Virchow1855] tends to obscure the deeper question of how the first cell came to be. Even within the origin-of-life community, research remains notably siloed: RNA-world chemists, mineral-surface and hydrothermal-vent advocates, astrobiologists, exoplanet scientists, and theoretical physicists working on stochastic thermodynamics, active matter, or phase transitions often pursue largely independent agendas. This fragmentation is widely recognised in the field, with repeated calls for integrating geochemistry, planetary science, and prebiotic chemistry [Sutherland2017EndBeginning, CleavesBada2012AlternativeNucleicAcids, Hazen2007emergence].
Against this backdrop of conceptual and disciplinary constraints, one thing is certain: with the rise of powerful AI tools and mechanistic models, we now have entirely new ways of exploring biological complexity. Tools like AlphaFold for protein folding [Jumper2021] and comprehensive whole-cell models [Crick1973_ProjectK, Karr2012, AhnHorst2022] allow us to estimate the information content of life using Kolmogorov (algorithmic) complexity. This measures the shortest description required to reproduce a system—the more structured or regular a system, the more it can be compressed and the easier it can be made [LiVitanyi2008]. Here, by combining such complexity estimates with basic rate-distortion theory [Shannon1959, CoverThomas2006], we aim to shed new light on the plausibility of life’s rapid emergence.
@@
We approach the problem of life’s origin by estimating how plausible it is for chemical processes to spontaneously assemble a sufficiently complex protocell within the available time on early Earth. We assume this window to be about $500$ Myr—interpolated between “optimistic” ($304$ Myr) and “conservative” ($939$ Myr) estimates [Kipping2020]. We apply a heuristic version of Shannon’s rate-distortion theory [Shannon1959, CoverThomas2006], which quantifies the minimal information rate $R(D)$ needed to construct a functioning system, given a tolerated distortion level $D$. Here, $D$ characterizes how much loss (error or degradation) can be tolerated in forming a viable protocell. The distortion rate depends on the entropy of the chaotic prebiotic environment, denoted $H_{\text{prebiotic}}$, which represents the information content of the chemical soup in a pond or hydrothermal vent.
@@
In this simplified formulation, we approximate $R(D)$, typically convex and decreasing, by the ratio of prebiotic entropy, $H_{\text{prebiotic}}$, and molecular persistence time, $D$, where $\eta<\!\!<1$ is an efficiency prefactor, introduced here as a compact way to account for microscopic inefficiencies (sometimes colloquially referred to as a “fudge factor”). The factor $\eta$, as harmless as it looks, accounts for the fact that only very few molecular encounters actually result in usable contributions. Molecules may be structurally misaligned, chemically unstable, or fail to integrate into a coherent assembly due to the absence of repair mechanisms or spatial constraints in the local environment. Determining this factor is really at the heart of the missing physical theory of protocell formation.
As one concrete illustration of the complexities subsumed into $\eta$, we consider the so-called homochirality problem, i.e. the difficulty of assembling a protocell in a racemic environment. Most biological molecules — amino acids, sugars, nucleotides — exist in two mirror-image forms (L and D), identical in energy but incompatible in function, yet all amino acids in biology are L and all ribose sugars in biology are D, requiring demixing and purification, effectively at a global scale. And even though the chiral symmetry breaking problem looks solved since the 1950s in a closed, well-mixed, kinetically idealized environment [Frank1953], it is hard to underestimate the problem of the constant influx of wrong chirality molecules from the outside, terminating biochemical pathways and thus jamming protocell assembly [Soai1995, Ozturk2023, Laurent2024].
Taken together, all quantities in Eq. 1—except perhaps $T_{\text{available}}$—are highly uncertain. Nevertheless, following this line of reasoning, the spontaneous origin of life remains plausible if the left-hand side exceeds the right-hand side. If not, we may be missing some crucial ingredient—such as additional evolutionary pathways, unknown physical principles, or external interventions like Crick and Orgel’s directed panspermia (or, more colorfully, aliens terraforming the planet). For the sake of argument, we set aside the ”cheap” statistical fallback that rare events inevitably occur somewhere in a vast universe.
Figure 2: Mathematical framework based on rate-distortion theory. (A) Funnel of collapsing entropy for the emergence of a protocell. While this is easily drawn, the challenge is to do this thermodynamically consistently. (B) Rate-distortion function from information theory, comparing $R(D)\sim H_{\text{prebiotic}}/D$ as a function of $D$ with $R_{\text{min}}$ from observational constraints (see the main text for explanations).
@@
In the absence of a time machine, we estimate conditions on Earth roughly 4 Gy ago. We begin with the left-hand side of Eq. 1, aiming to estimate the entropy of the chemical soup, $H_{\text{prebiotic}}$, and the tolerated error rate DD—which characterizes how long complex organic molecules persist before degrading. Importantly, the entropy $H_{\text{prebiotic}}$ pertains not to an entire ocean or vent system, but to the volume accessible to a nascent protocell. While the volume of a single protocell may be around 1 femtolitre ($10^{-15}$ l), relevant molecules could diffuse in from a larger surrounding droplet. We seek order-of-magnitude estimates informed by chemical experiments simulating early Earth conditions, meteorite data, and planetary spectroscopy.
@@
In the Miller–Urey experiments, spark discharges in a reducing gas mixture produced a variety of amino acids and other small organic molecules [Miller1953]. During the Cassini–Huygens mission, analysis of Titan’s atmosphere revealed that complex photochemistry can generate thousands of distinct organic species, including nitriles, hydrocarbons, and possible prebiotic precursors [Horst2017]. Similarly, carbonaceous chondrite meteorites (e.g., Murchison) contain more than 60 amino acids, as well as nucleobases, polyols, carboxylic acids, and related compounds [Sephton2002]. This is consistent with samples returned from asteroid Bennu by the OSIRIS-REx mission, which revealed a rich inventory of organic compounds—including those preserved in salt-rich brine inclusions in phyllosilicate minerals, dating back to $\sim 4.5$ billion years [McCoy2025].
@@
Combinatorial chemistry across diverse geochemical settings could plausibly yield thousands of unique organic molecules, suggesting a chemical library on the order of $10^{5}$$10^{6}$ small molecules in the prebiotic environment. Assuming an average information content of $\sim 10$ bits per molecule—representing functional or structural specificity—this allows for roughly $10^{3}$ distinguishable states per molecule (e.g., different side chains, folding propensities, or reactivities). This estimate aligns with protein domain annotations and bits-per-residue values from sequence–function maps [Durston2007, AdamiCerf2000, Weiss2000ProteinInfo], including quantitative analyses of sequence information content in natural proteins [Possenti2018ProteinInfoPartition]. Related bounds also arise from canonical chemical representations such as SMARTS and InChI, whose string lengths provide effective upper limits on molecular descriptive complexity [Heller2015]. Multiplying, we estimate the prebiotic entropy as $H_{\text{prebiotic}}\sim 10^{6}$$10^{7}$ bits.
@@
Next, we consider the distortion timescale $D$, interpreted as the average lifetime of a complex organic molecule before degradation by hydrolysis, UV damage, oxidation, or further transformation. This defines the “error rate” or persistence window for useful molecular structures. Under plausible early-Earth conditions, complex organics such as ribose and peptides have half-lives ranging from hours to days. For example, ribose degrades with a half-life of $\sim 3$ hours at pH 7 and 37C [Larralde1995]. Small peptides degrade over hours to days depending on sequence and pH. Nucleobases like adenine are sensitive to solar UV and degrade within days to weeks. Without an ozone layer, degradation could be faster—on the scale of hours to days. Conservatively, we adopt a mean lifetime of $\sim 1$ day, or about $10^{5}$ seconds, for moderately stable molecules in a warm, aqueous, anoxic environment exposed to UV. However, in cryogenic or mineral-protected contexts—such as asteroid interiors—organic molecules may persist for billions of years [McCoy2025]. The existence of such intergalactic freezers implies that $D$ could vary over many orders of magnitude depending on local geochemical context.
@@
If the molecular lifetime were much shorter—say $D\sim 1$ s, due to rapid diffusion or lack of protective mechanisms—then the required information rate would increase accordingly. The key idea remains: if $R(D)$ is too low to meet $R_{\text{min}}=I_{\text{protocell}}/T_{\text{available}}$, then abiotic emergence becomes implausible without invoking directed evolution, external contributions (e.g., panspermia), or unknown self-organizing principles. The delivery of relatively stable, complex molecules—by meteorites or interstellar dust—could act as a bootstrap for prebiotic evolution [McCoy2025]. Given the large uncertainties in parameter values, our aim is not to deliver a definitive conclusion but to structure a quantitative discussion.
@@
The right-hand side of Eq. 1 requires an estimate of the complexity of a protocell, $I_{\text{protocell}}$. How can we quantify this complexity, including its structural and dynamic organization? It shouldn’t come as a surprise that cells are not simply bags of well-mixed molecules [deDuve1984GuidedTour], but highly structured systems—down to the folding and interaction of their constituent proteins. Moreover, cellular processes are orchestrated in time. If proteins are the musicians, then signaling, metabolic, and genetic pathways are the conductor and the score — guiding timing, harmony, and transitions.
@@
where $I_{g}$ is the genetic information, $I_{s}$ the structural information (e.g. protein folds, cellular organization), and $I_{d}$ the dynamic information encoded in biochemical pathways. Each term is estimated via Kolmogorov (algorithmic) complexity, defined as the length of the shortest program that can realistically generate the relevant component, at least conceptually. Hence, Kolmogorov complexity describes the information needed to specify the final structure.
However, this description is not yet sufficient. Even if all components were fully specified, physical realization introduces an additional and independent constraint: the complexity of assembly, which captures not only the number of components but also the order, feasibility, and reliability of the assembly steps themselves. While Kolmogorov complexity quantifies the length of a minimal description, assembly complexity reflects the execution of that description under physical constraints. We therefore introduce an additional contribution, $I_{a}$, accounting for assembly in Eq. 3. This term emphasizes that random assembly cannot generate a complex cell, because the dominant informational bottleneck lies not in specification but in the assembly pathway. Viable biological organization requires nested, hierarchical construction: elementary components assemble into intermediate modules, which in turn form higher-level structures, culminating in the protocell. Each stage reduces entropy locally, but at the cost of increasing dependence on correct ordering and stabilization. Consequently, early protocells must have been extremely simple, with complexity growing only incrementally, as each new organizational layer required robust scaffolding to stabilize those below.